OpenStudy (aaronrocky):

Puzzle question, fairly easy My answer on this is - not enough information to solve algebraically, you have to do it by trial and error. Am I wrong? "Difference between 2 numbers is 3. 4x the smaller is divided by the larger, the quotient is 5. Find numbers"

OpenStudy (johnweldon1993):

Hmm *on my phone so latex won't work too well but...* Let "x" be the smaller number And "y" be the larger number x-y=3 (4x)/(y) = 5 Let's solve the top equation for "x" x=3+y Plug that into the second equation and solve for "y"

OpenStudy (aaronrocky):

I'm confused, if X and Y are givens why are they set up in that equation? How would I "solve for X" when X is any number I pick?

OpenStudy (triciaal):

X and Y are not given. you have the relationship between them so that for any number you pick and you know the relationship then you can find Y Here the question is really asking what numbers were picked

OpenStudy (phi):

In general, if you have two unknowns, then you need two equations If you have two equations, you (usually) have enough info.

OpenStudy (phi):

In your problem you have two equations: "Difference between 2 numbers is 3: x-y= 3 (we assume x is bigger than y) 4 times the smaller is divided by the larger, the quotient is 5. the smaller is (assumed) to be y. 4 times the smaller is 4y divided by the larger (x): 4y/x= 5 now you use algebra (if you know how) for example, using 4y/x= 5 multiply both sides by x: 4y= 5x now divide both sides by 4: y= (5/4)x use that "value" for y in the first equation x - y = 3 becomes \[ x-\frac{5}{4}x= 3\]